200 =600(a+1)/24
200 x 24 = 600(a+1) (by multiplying both sides of the equation by 24)
4800 = 600(a+1) (by simplifying the left hand side)
4800/600 = a+1 (by dividing both sides of the equation by 600)
8 = a+1 (by simplifying the left hand side)
8-1 =a (subtracting 1 from both sides of the eqauiton)
7 =a (simplifyng the left hand side)
Therefore
a =7




Answer by ainsworth50(1)

200=600(a+1)/24
200 X 24 = 600(a+1) (I multiplied both sides of the equation by 24)
4800 = 600(a+1) ( I simplified the left hand side)
4800/600 = a+1 (By dividing both sides by 600)
8 = a+1 (Again by simplifying the left)
8-1 = a (I subtracxtd 1 from both sides)
7 = a ( Working out 7-1)
Therefore
a =7 ( The last step is not really needed. It says the same
thing as the one before)